OpenStudy (hba):

Which order of differentiation will calculate f(x,y) faster: x first or y first? Try to answer without writing anything down. f(x,y)=xsiny+e^y

OpenStudy (hba):

@hartnn

hartnn (hartnn):

what do u need to calculate ?

hartnn (hartnn):

f'(x,y) ?

OpenStudy (hba):

I saw this question in mixed partial derivatives section It's just asking fxy(x,y) would be faster or fyx(x,y)

OpenStudy (jhannybean):

\[\begin{align} f(x,y) = x\sin(y)+e^y & \\ &: f_x(x,y) = \sin(y) +0 \\ &:f_{xy}(x,y) = \cos(y) \\ \\ &:f_y(x,y) = x\cos(y)+e^y \\ &:f_{yx}(x,y) = \cos(y) \\ \end{align}\] This si Clairaut's Theorem, isn't it?

OpenStudy (anonymous):

I would say \(x\) first.

OpenStudy (anonymous):

me goin with y

OpenStudy (hba):

@wio Is it just a random guess or is their a way to do it? I know how to solve it @jhanny

OpenStudy (anonymous):

Well, \(at+b\) is much more simple than \(a\sin t+e^t\).